OpenStudy (anonymous):

If c > 0, prove that (ca,cb) =c(a,b).

OpenStudy (aravindg):

well geometrically speaking if (a,b) is a point and we multiply it by c ,then both a and b will get multiplied by c. lets see if i can provide an algebraic proof

OpenStudy (anonymous):

i think (a,b) represents the GCD

OpenStudy (anonymous):

Let a=nx and b=ny where n is the GCD of a and b then, ca=c*a=c*n*x [ NOTE c>0] cb=c*b=c*n*y Thus, (ca,cb)=cn ALSO, c(a,b)=c*n=cn Thus, If c > 0, (ca,cb) =c(a,b)

OpenStudy (anonymous):

Thank you for your help today.

OpenStudy (anonymous):

welcome